You would like to set aside enough money to pay for the maintenance you will need for your new car. You estimated that you will need $872 at the end of year 2, and another $1407 at the end of year 7. How much do you need to set aside today in order to pay for all of these expenses, assuming that the bank pays 8% per year compounded annually? Enter your answer as follows: 1234 Round your answer. Do not use a dollar sign ($), a comma () or any decimal points()
Answer #1

Solution: 1569

Working:

1) Given: FV = $872; n =2; i = 8%

PV = ?

FV = PV (1 + R/100)^n

872 = PV ( 1 + 0.08)^2

PV = 872 / 1.1664 = 747.6

2) Given: FV = $1407; n =7; i = 8%

PV = ?

FV = PV (1 + R/100)^n

1407 = PV ( 1 + 0.08)^7

PV = 1407 / 1.7138 = 820.97

Thus total amount to be saved today = 747.6 + 820.97 = 1569